How would I graph y=x^2-10
I already know my vertex is (0,-10) so I would start with these points but where do I put my other points?
The quiz I'm using has be placing 2 points one of which being (0,-10) I just dont know how to find the next....

Answer by TimothyLamb(4379)

y = x^2 - 10
---
the above quadratic equation is in standard form, with a=1, b=0, and c=-10
---
to solve the quadratic equation, by using the quadratic formula, plug this:
1 0 -10
into this: https://sooeet.com/math/quadratic-equation-solver.php
---
the two real roots (i.e. the two x-intercepts), of the quadratic are:
x = 3.16227766, y = 0
x = -3.16227766, y = 0
---
the quadratic vertex is a minimum at: ( 0, -10 )
---
now you have 3 different points to plot.
if you want more point, just calculate them like so:
x = 1, y = (1)^2 - 10 = -9
x = 2, y = (2)^2 - 10 = -6
x = 3, y = (3)^2 - 10 = -1
---
the axis of symmetry is: x = 0
so you can reflect the above x,y points about this axis of symmetry to get the corresponding points for x=-1, x=-2, x=-3, etc.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php